OpenStudy (anonymous):

Can anyone help me with this question please? Y=cos^-1(sin x+5).

OpenStudy (anonymous):

yeah, forget it. the input is between 4 and 6 and the domain of arccosine is \([0,\pi]\)

OpenStudy (anonymous):

you cannot take the inverse cosine of a number larger than \(\pi\) so this does not make sense

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

in other words \[\cos^{-1}(\sin(x)+5)\] does not exist what was the actual question?

OpenStudy (anonymous):

it looks like this \[Y=\cos^{-1} (\sin x+5)\]

OpenStudy (anonymous):

derive this using the chain rule \[\frac{ d }{ dx }(\cos^{-1}(\sin(x)+5)=\frac{ -\frac{ du }{ dx } }{ \sqrt{1-u^{2}} }\] where \(u=\sin(x)+5\) and \(\frac{d cos^{-1}(u) }{du}=-\frac{ 1}{ \sqrt{1-u^{2}} }\)

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

it gives you \[-\frac{ \frac{ d }{ dx } (\sin(x)+5)}{ \sqrt{1-(\sin(x)+5)^{2}} }\]

OpenStudy (anonymous):

then derive each term on the top, then you will get your answer.

OpenStudy (anonymous):

can you please help me all the to the answer then I ll work with my other problems too

OpenStudy (anonymous):

So this way I ll know the right method to do that

OpenStudy (anonymous):

if you dont mind sir?

OpenStudy (anonymous):

do you know how to derive sin(x)? and derive 5? because the bottom stays the same and the only thing you need to do is derive sin(x) + 5 on the top

OpenStudy (anonymous):

yeah kind of but not sure professor did go over but the way she taught was really confusing

OpenStudy (anonymous):

you know the derivative of any number with respect to an unknown, in this case, x, is 0 right? and then can you derive sin(x) on your own?

OpenStudy (anonymous):

hmmm i know that but just to make it sure can you please go over with me to the answer?

OpenStudy (anonymous):

the derivative of \(sin (x) \) is \(cos (x) \)

OpenStudy (anonymous):

yup

OpenStudy (anonymous):

then you have your answer. :)